# Strategic Tradeoffs Between Humans and AI in Multi-Agent Bargaining

### v3

@@ A.3 Multi-agent trading algorithm with Bayesian learning
[…]
@@
      $\max_{(\mathbf{x},\mathbf{y})}\sum_{\mathbf{v}_{-i}}\mathbf{1}\Bigl\{\exists\,\text{receiver }r\neq i:\textsc{accept}\bigl(v_{r};\mathbf{x},\mathbf{y}\bigr)\Bigr\}\times\Bigl(u_{i}\bigl(h_{i}-\mathbf{x}+\mathbf{y}\bigr)-u_{i}\bigl(h_{i}\bigr)\Bigr)\times B_{i}(\mathbf{v}_{-i}),$
@@ Appendix B Human data collection
B.1 Implementation on Deliberate Lab
Platform.
The bargaining game was implemented via Deliberate Lab (Tsai et al. (2024)), an open-source platform for conducting group human-LLM social science experiments. We will publish the bargaining game implementation within the platform upon acceptance.
@@ B.1 Implementation on Deliberate Lab
Upon entering Deliberate Lab through a web link, participants proceed through a multi-stage experiment including informed consent, game instructions, comprehension checks, and payout information. Upon completing the final comprehension check, they wait in a “Lobby” stage for other participants. When three participants are in the lobby, they are sent an invitation to join a live bargaining game (Fig 6). Following the game, there is a post-game survey. For anonymity, we used a Deliberate Lab feature that assigns participants an anonymous animal avatar (e.g., “Bear”) as they join the experiment.
Figure 6: The game interface shown to participants. This shows a participant’s game state. The top left indicates that they are currently the Proposer, offering an interface to propose an offer. A table below displays all players’ chip volumes. To the right, the public ledger shows a history of game rounds, turns, and trade statuses.
@@
To submit trade proposals, participants used text and selection fields to specify number and type of chip, respectively. The interface automatically calculated and displayed projected total chip value if the proposal were to be accepted (Fig 6); it also prevented players from submitting invalid offers by disabling the "submit" button.
For accepting or rejecting proposals, the interface calculated and displayed the projected total chip value if the player were to accept the proposal. It also disabled acceptance for offers the player could not fulfill (Fig 7, bottom). All participants were shown a table of all players’ chip quantities and a log of actions (i.e., when trades were proposed, accepted, and rejected, and when each round and turn of the game changed). Turn order was determined randomly.
Figure 7: Accepting or rejecting a trade. The bottom image shows a disabled “accept” option.
@@
To conduct experiments, the experimenter used Deliberate Lab’s dashboard (Fig 8) to monitor participants as they progressed through the experiment’s information and comprehension check stages, then manually transferred participants into groups of three to play the bargaining game once the player was ready to be transferred. The experimenter was also able to send attention checks to participants (e.g., if several minutes elapsed without any visible progress) and remove them from the experiment (e.g., if attention checks went unanswered).
Figure 8: Experiment dashboard for managing participants. On the left, experimenters can view participants as they enter the experiment, send attention checks, and/or remove participants. On the right, experimenters can monitor the status of a selected participant and assign them to a cohort of three for gameplay.
@@ Appendix C Post-game participant survey
C.1 Survey questions
Table 1: Post-game survey questions.
@@ C.1 Survey questions
3 On a scale from 1 to 10, how would you rate your trading strategy, where 1 is highly competitive (focused mainly on your own gains) and 10 is highly collaborative (focused on other players’s potential gains and mutual benefits? 1 through 10 scale; 1: "Not at all aggressive", 10: "Very aggressive"
4 On a scale from 1 to 10, how certain are you that you made the best possible trades during the experiment? 1 through 10 scale; 1: "Not at all confident", 10: "Very confident"
5 On a scale from 1 to 10, how satisfied are you with your final trading outcomes? 1 through 10 scale; 1: "Not at all satisfied", 10: "Very satisfied"
6 On a scale of 1 to 10, how would you rate the mental effort you put into today’s trading games, where 1 means ’I barely engaged in any thinking’ and 10 means ’I put in significant mental effort’? 1 through 10 scale; 1: "No mental exertion at all", 10: "Very high mental exertion"
@@
C.2 Quantitative survey responses
Table 2: Summary statistics of self-assessment variables (mean and standard error). There was no significant change in certainty or exertion as game complexity increased.
Game variation Competitive Certainty Satisfaction Exertion
2 chip 5.57 (0.34) 5.57 (0.33) 6.68 (0.31) 7.11 (0.24)
3 chip 5.32 (0.34) 5.92 (0.33) 7.07 (0.30) 7.10 (0.25)
4 chip 5.75 (0.34) 6.04 (0.33) 7.46 (0.26) 7.46 (0.19)
@@ Appendix D Additional clarification on strategic regret classifications
Figure 9: Regret behavior visualization: The top-left panel shows an overview trajectory in Fig 4. The top-right panel illustrates a “No Regret” scenario for P2 as a proposer. The bottom-left panel depicts “Forced Regret” for P3 as an acceptor. The bottom-right panel shows “Unforced Regret” for P2 as a decliner at Turn 8. Solid lines represent actual player paths, while dashed lines indicate counterfactual paths.
@@
Here, we utilized counterfactual analysis to visually dissect the three kinds of regret behavior previously discussed in Section 6.2.2, as illustrated in Fig  9. In these visualizations, the player’s actual path of total chip value is depicted by a solid colored line, while the counterfactual path—representing the outcome had a different decision been made — is indicated by a gray dashed line.
Counterfactual definition:
@@
The counterfactual path is constructed by simulating an alternative decision at a critical juncture (e.g., accepting a declined offer, declining an accepted offer, or making a different proposal) and projecting the subsequent value trajectory. This allows for a direct comparison between the actual outcome and what could have been.
No regret scenarios.
@@
This scenario is demonstrated by Player 2 (P2) in the role of a proposer, as shown in the upper-right panel of Fig 9, focusing on Turn 9. Player 2 makes a proposal that is accepted, leading to an actual chip value of $15.3$. The counterfactual path, representing a scenario where P2 might not have made this proposal or made a less optimal one, results in a lower value of 14.9. In this “No Regret” instance, the player’s actual total value path is superior to or equal to the evaluated counterfactual path, indicating that the decision made was optimal given the available information and subsequent events.
Forced regret scenarios.
@@
The lower-left panel of Fig 9 illustrates “Forced Regret” experienced by Player 3 (P3) as an acceptor from Turn 1 to Turn 2. P3 initially accepts an offer from P1 at Turn 1, resulting in an actual value of 10.8. However, a more lucrative offer from P2 becomes available in Turn 2. Due to the commitment made in Turn 1 (e.g., changed chip inventory), P3 is unable to capitalize on this subsequent, better opportunity. The counterfactual path shows that had P3 declined P1’s initial offer, they could have potentially accepted P2’s offer, leading to a higher value of 11.0.
Unforced regret scenarios.
@@
The lower-right panel of  9 demonstrates Player 2 (P2) in the role of a decliner from Turn 8 to 9. P2 is presented with a profitable trade offer but chooses to decline it. P2’s actual chip value remains at 14.9 after this decision. The counterfactual path, however, demonstrates that had P2 accepted this profitable offer, their chip value would have increased to 15.0. This highlights an avoidable error: a superior option was available and actionable, but the agent failed to select it, leading to a sub-optimal outcome. Unlike forced regret, no prior decision prevented P2 from taking the better option; the regret stems directly from the choice made at that specific juncture.
@@ Appendix E Surplus values and significance
Table 3: Scaled final surplus gain (mean and standard error). This is the ratio of observed total surplus to optimal surplus.
Game variation Human GPT-4o GPT-4o refined Gemini Gemini refined Bayesian agent
2 chip 0.60 (0.06) 0.69 (0.03) 0.68 (0.04) 0.42 (0.04) 0.45 (0.05) 0.74 (0.04)
3 chip 0.59 (0.04) 0.62 (0.03) 0.64 (0.03) 0.42 (0.03) 0.43 (0.03) 0.80 (0.03)
4 chip 0.54 (0.03) 0.54 (0.02) 0.58 (0.02) 0.37 (0.02) 0.33 (0.02) 0.73 (0.02)
Table 4: Two-sided t-test p-values comparing surplus values across populations (N=144 per game). Arrows indicate whether the column’s mean surplus was higher (↑) or lower (↓) than the row’s.
@@
GPT-4o 0.302 ↓ 0.000 ↓ 0.000 ↓ 0.141 ↑
GPT-4o refined 0.000 ↓ 0.000 ↓ 0.414 ↑
Gemini 0.359 ↑ 0.000 ↑
Gemini refined 0.000 ↑
@@
3 chip
Human 0.525 ↑ 0.021 ↑ 0.001 ↓ 0.000 ↓ 0.000 ↑
GPT-4o 0.164 ↑ 0.000 ↓ 0.004 ↓ 0.000 ↑
GPT-4o refined 0.000 ↓ 0.000 ↓ 0.001 ↑
Gemini 0.799 ↑ 0.000 ↑
Gemini refined 0.000 ↑
@@
4 chip
Human 0.601 0.258 ↑ 0.000 ↓ 0.000 ↓ 0.000 ↑
GPT-4o 0.270 ↑ 0.000 ↓ 0.000 ↓ 0.000 ↑
GPT-4o refined 0.000 ↓ 0.000 ↓ 0.000 ↑
Gemini 0.003 ↓ 0.000 ↑
Gemini refined 0.000 ↑
@@
Appendix F Aside on smaller model performance
We tested smaller models from both the Gemini and GPT families— Gemini-2.5-Flash and GPT-o4-mini—and found that their overall performance falls short of both human participants and larger models such as GPT-4o, particularly in terms of surplus generation. Even with prompt refinement, the average surplus gain remains substantially lower.
Procedurally, these small models also behave quite differently from their larger counterparts. As shown in Figure 10, proposals by Gemini-2.5-Flash and GPT-4-mini are heavily concentrated around a narrow region: nearly zero net surplus gain and a 1:1 trade ratio. Roughly 90% of offers consist of exchanging one chip for exactly one of another type (e.g., 1 red for 1 green). Loss-incurring proposals are virtually nonexistent—but so are high-surplus trades (e.g., net surplus >2>2). Prompt refinement slightly broadens the distribution but fails to shift the models away from this dominant 1:1 pattern.
Figure 10: Comparison of the trade space (3-Chip) between small LLM agents (GPT-o4-mini and Gemini-2.5-flash).
F.1 Full strategic analysis visualization
Figure 11: This visualization follows the same structure as Figure 5, with additional Refined LLMs. Strategic differences between refined and out-of-the-box LLMs are empirically similar.
F.2 Supplementary trade space visualizations
Summary statistics of trade space values are in Table 5, with the 2-chip visualization in Figure 12 and 4-chip visualization in Figure 13. They exhibit similar behaviors as discussed in Section 6.2.1.
Figure 12: Comparison of the trade space (2-Chip) between Human participants, LLM agents (GPT-4o and Gemini-1.5 Pro), and Bayesian agents.
Figure 13: Trade space comparison (4-chip) between humans, LLM agents (GPT-4o and Gemini-1.5 Pro), and Bayesian agents.
Table 5: Summary statistics from trade space visualizations.
@@ F.2 Supplementary trade space visualizations
Appendix G LLM simulation prompts
G.1 Rule explanation
@@
The game consists of 3 rounds of trading. During each round, each player will have a turn to propose 1 trade. These turns are pre-determined in a random order and the order stays the same in each round.
@@
G.2 Proposer’s prompt
We used a Chain-of-Thought approach to elicit LLM agent’s trade proposals:

### v4

@@ A.3. Multi-Agent Trading Algorithm with Bayesian Learning
2:   $\mathbf{I}$: set of players
3:   $\mathbf{G}$: set of goods
4:   $\mathbf{A^{0}}$: initial allocation $\{a_{ig}\}$ for all $i\in\mathbf{I},g\in\mathbf{G}$
5:   $\{B_{i}(\mathbf{v}_{-i})\}_{i\in\mathbf{I}}$: each player’s prior belief over other players’ valuations
@@
      (3) $\max_{(\mathbf{x},\mathbf{y})}$ $\sum_{\mathbf{v}_{-i}}\mathbf{1}\!\left\{\exists r\neq i:\ \textsc{accept}\bigl(v_{r};\mathbf{x},\mathbf{y}\bigr)\right\}$
      $\qquad\qquad\times\bigl(u_{i}(h_{i}-\mathbf{x}+\mathbf{y})-u_{i}(h_{i})\bigr)\times B_{i}(\mathbf{v}_{-i}).$
@@ Appendix B Human Data Collection
B.1. Implementation on Deliberate Lab
@@
Upon entering Deliberate Lab through a web link, participants proceed through a multi-stage experiment including informed consent, game instructions, comprehension checks, and payout information. Upon completing the final comprehension check, they wait in a “Lobby” stage for other participants. When three participants are in the lobby, they are sent an invitation to join a live bargaining game (Figure 6). Following the game, there is a post-game survey. For anonymity, we used a Deliberate Lab feature that assigns participants an anonymous animal avatar (e.g., “Bear”) as they join the experiment.
Figure 6. The game interface exactly as it was shown to participants. The top left indicates that they are currently the proposer, offering an interface to propose an offer. A table below displays all players’ chip volumes. To the right, the public ledger shows a history of game rounds, turns, and trade statuses.
The game interface exactly as it was shown to participants. The top left indicates that they are currently the proposer, offering an interface to propose an offer. A table below displays all players’ chip volumes. To the right, the public ledger shows a history of game rounds, turns, and trade statuses.
Figure 7. Experiment dashboard for managing participants. On the left, experimenters can view participants as they enter the experiment, send attention checks, and/or remove participants. On the right, experimenters can monitor the status of a selected participant and assign them to a cohort of three for gameplay.
Experiment dashboard for managing participants. On the left, experimenters can view participants as they enter the experiment, send attention checks, and/or remove participants. On the right, experimenters can monitor the status of a selected participant and assign them to a cohort of three for gameplay.
@@
To submit trade proposals, participants used text and selection fields to specify the number and type of chip, respectively. The interface automatically calculated and displayed the projected total chip value if the proposal were to be accepted (Figure 6); it also prevented players from submitting invalid offers by disabling the “submit” button. For accepting or rejecting proposals, the interface calculated and displayed the projected total chip value if the player were to accept the proposal. It also disabled acceptance for offers the player could not fulfill (Figure 8, bottom). All participants were shown a table of all players’ chip quantities and a log of actions (i.e., when trades were proposed, accepted, and rejected, and when each round and turn of the game changed). Turn order was determined randomly.
@@
To conduct experiments, the experimenter used Deliberate Lab’s dashboard (Figure 7) to monitor participants as they progressed through the experiment’s information and comprehension check stages, then manually transferred participants into groups of three to play the bargaining game once the player was ready to be transferred. The experimenter was also able to send attention checks to participants (e.g., if several minutes elapsed without any visible progress) and remove them from the experiment (e.g., if attention checks went unanswered).
Figure 8. Accepting or rejecting a trade. The bottom image shows a disabled “accept” option.
Accepting or rejecting a trade. The bottom image shows a disabled “accept” option.
@@ Appendix C Post-Game Participant Survey
Game variation Competitive Certainty Satisfaction Exertion
2-chip 5.57 (0.34) 5.57 (0.33) 6.68 (0.31) 7.11 (0.24)
3-chip 5.32 (0.34) 5.92 (0.33) 7.07 (0.30) 7.10 (0.25)
4-chip 5.75 (0.34) 6.04 (0.33) 7.46 (0.26) 7.46 (0.19)
Table 1. Summary statistics of self-assessment Likert variables (mean and standard error); full survey questions in Table 2. There was no significant change in certainty or exertion as game complexity increased.
The survey questions are provided in Table 2. Responses to Likert-scale questions are aggregated in Table 1. Qualitative thematic analysis is provided in Figure 9.
Table 2. Post-game survey questions.
@@
3 On a scale from 1 to 10, how would you rate your trading strategy, where 1 is highly competitive (focused mainly on your own gains) and 10 is highly collaborative (focused on other players’ potential gains and mutual benefits) 1 through 10 scale; 1: “highly competitive”, 10: “highly collaborative”
4 On a scale from 1 to 10, how certain are you that you made the best possible trades during the experiment? 1 through 10 scale; 1: “Not at all confident”, 10: “Very confident”
5 On a scale from 1 to 10, how satisfied are you with your final trading outcomes? 1 through 10 scale; 1: “Not at all satisfied”, 10: “Very satisfied”
6 On a scale of 1 to 10, how would you rate the mental effort you put into today’s trading games, where 1 means ’I barely engaged in any thinking’ and 10 means ’I put in significant mental effort’? 1 through 10 scale; 1: “No mental exertion at all”, 10: “Very high mental exertion”
@@
Figure 9. Distribution of self-reported strategies across game complexity (thematic analysis). Participants described their trading approach as focused on profit maximization, collaborative play, or color optimization. Profit-focused strategies were most common (112 responses), while collaborative (56) and color-focused (55) strategies appeared at similar rates. The distribution across 2-, 3-, and 4-chip games was relatively consistent within each strategy type.
Distribution of self-reported strategies across game complexity (thematic analysis). Participants described their trading approach as focused on profit maximization, collaborative play, or color optimization. Profit-focused strategies were most common (112 responses), while collaborative (56) and color-focused (55) strategies appeared at similar rates. The distribution across 2-, 3-, and 4-chip games was relatively consistent within each strategy type.
@@ Appendix D Additional Clarification on Strategic Regret Classifications
Figure 10. Regret behavior visualization: The top-left panel shows an overview trajectory from Figure 4. The top-right panel illustrates a “No Regret” scenario for P2 as a proposer. The bottom-left panel depicts “Forced Regret” for P3 as an acceptor. The bottom-right panel shows “Unforced Regret” for P2 as a decliner at Turn 8. Solid lines represent actual player paths, while dashed lines indicate counterfactual paths.
[…]
@@
Here, we utilized counterfactual analysis to visually dissect the three kinds of regret behavior previously discussed in Section 6.2.2, as illustrated in Figure 10. In these visualizations, the player’s actual path of total chip value is depicted by a solid colored line, while the counterfactual path—representing the outcome had a different decision been made—is indicated by a gray dashed line.
@@
Counterfactual definition: The counterfactual path is constructed by simulating an alternative decision at a critical juncture (e.g., accepting a declined offer, declining an accepted offer, or making a different proposal) and projecting the subsequent value trajectory. This allows for a direct comparison between the actual outcome and what could have been.
@@
No regret scenarios. This scenario is demonstrated by Player 2 (P2) in the role of a proposer, as shown in the upper-right panel of Figure 10, focusing on Turn 9. Player 2 makes a proposal that is accepted, leading to an actual chip value of $15.3$. The counterfactual path, representing a scenario where P2 might not have made this proposal or made a less optimal one, results in a lower value of 14.9. In this “No Regret” instance, the player’s actual total value path is superior to or equal to the evaluated counterfactual path, indicating that the decision made was optimal given the available information and subsequent events.
@@
Forced regret scenarios. The lower-left panel of Figure 10 illustrates “Forced Regret” experienced by Player 3 (P3) as an acceptor from Turn 1 to Turn 2. P3 initially accepts an offer from P1 at Turn 1, resulting in an actual value of 10.8. However, a more lucrative offer from P2 becomes available in Turn 2. Due to the commitment made in Turn 1 (e.g., changed chip inventory), P3 is unable to capitalize on this subsequent, better opportunity. The counterfactual path shows that had P3 declined P1’s initial offer, they could have potentially accepted P2’s offer, leading to a higher value of 11.0.
@@
Unforced regret scenarios. The lower-right panel of Figure 10 demonstrates Player 2 (P2) in the role of a decliner from Turn 8 to 9. P2 is presented with a profitable trade offer but chooses to decline it. P2’s actual chip value remains at 14.9 after this decision. The counterfactual path, however, demonstrates that had P2 accepted this profitable offer, their chip value would have increased to 15.0. This highlights an avoidable error: a superior option was available and actionable, but the agent failed to select it, leading to a sub-optimal outcome. Unlike forced regret, no prior decision prevented P2 from taking the better option; the regret stems directly from the choice made at that specific juncture.
@@ Appendix E Surplus Values and Significance
Surplus values are provided in Table 3 and Table 4.
Population 2-chip 3-chip 4-chip
Human 0.60 (0.06) 0.59 (0.04) 0.54 (0.03)
GPT-4o 0.69 (0.03) 0.62 (0.03) 0.54 (0.02)
GPT-4o refined 0.68 (0.04) 0.64 (0.03) 0.58 (0.02)
Gemini 0.42 (0.04) 0.42 (0.03) 0.37 (0.02)
Gemini refined 0.45 (0.05) 0.43 (0.03) 0.33 (0.02)
Bayesian agent 0.74 (0.04) 0.80 (0.03) 0.73 (0.02)
Table 3. This table provides the means and standard errors for the average ratio of observed total surplus to optimal surplus achieved across agent-game combinations. It corresponds to the numbers in Figure 2.
@@
GPT-4o 0.302 ↓ $p<0.001$ $p<0.001$ 0.141 ↑
GPT-4o refined $p<0.001$ $p<0.001$ 0.414 ↑
Gemini 0.359 ↑ $p<0.001$
Gemini refined $p<0.001$
@@
3-chip
Human 0.525 ↑ 0.021 ↑ 0.001 ↓ $p<0.001$ $p<0.001$
GPT-4o 0.164 ↑ $p<0.001$ 0.004 ↓ $p<0.001$
GPT-4o refined $p<0.001$ $p<0.001$ 0.001 ↑
Gemini 0.799 ↑ $p<0.001$
Gemini refined $p<0.001$
@@
4-chip
Human 0.601 0.258 ↑ $p<0.001$ $p<0.001$ $p<0.001$
GPT-4o 0.270 ↑ $p<0.001$ $p<0.001$ $p<0.001$
GPT-4o refined $p<0.001$ $p<0.001$ $p<0.001$
Gemini 0.003 ↓ $p<0.001$
Gemini refined $p<0.001$
@@
Table 4. Two-sided t-test p-values comparing surplus values across populations (144 games). Arrows indicate whether the column’s mean surplus was higher (↑) or lower (↓) than the row’s.
Appendix F Additional Data Analysis
F.1. Learning Effects
In Figure 11, we investigate learning effects for humans in the 2, 3, and 4-chip variants of the negotiation game by comparing the final surplus gain across the two stages of play. We see that while there is a small increase in performance in each game, it is not statistically significant. We also cannot determine how the complexity of the game relates to learning effects, as the 3-chip game seems to have the smallest effect.
Figure 11. Between stage (game) comparisons of human group surplus by game variant.
Between stage (game) comparisons of human group surplus by game variant.
F.2. Supplementary Trading Pattern Visualizations
Summary statistics of trading patterns are provided in Table 5, and average final surplus information in Table 3. The 2-chip visualization is shown in Figure 12 and 4-chip visualization in Figure 13. They exhibit similar behaviors as discussed in Section 6.2.1.
Figure 12. This figure is identical to Figure 3 in the main text, but provides the trading patterns for the 2-chip game.
[…]
Figure 13. This figure is identical to Figure 3 in the main text, but provides the trading patterns for the 4-chip game.
[…]
@@ F.2. Supplementary Trading Pattern Visualizations
Table 5. Summary statistics of the trading patterns in Figures 3, 12, and 13.
Appendix G Aside on Smaller Model Performance
We tested smaller models from both the Gemini and GPT families—Gemini-2.5-Flash and GPT-o4-mini—and found that their overall performance falls short of both human participants and larger models such as GPT-4o, particularly in terms of surplus generation. Even with prompt refinement, the average surplus remains substantially lower.
Procedurally, these small models also behave quite differently from their larger counterparts. As shown in Figure 14, proposals by Gemini-2.5-Flash and GPT-o4-mini are heavily concentrated around a narrow region: nearly zero net surplus gain and a 1:1 trade ratio. Roughly 90% of offers consist of exchanging one chip for exactly one of another type (e.g., 1 red for 1 green). Rarely do small LLMs propose trades that would result in a loss. However, they also rarely propose trades that would benefit them substantially (e.g., a net surplus of >2>2). Prompt refinement slightly broadens the distribution but fails to shift the models away from this dominant 1:1 pattern.
Figure 14. This figure is identical to Figure 3 in the main text, but provides the trading patterns for the smaller models
[…]
Appendix H LLM Simulation Prompts
H.1. Rule Explanation
@@
The game consists of 3 rounds of trading. During each round, each player will have a turn to propose 1 trade. These turns are pre-determined in a random order, and the order stays the same in each round.
@@
H.2. Proposer’s Prompt